OpenStudy (anonymous):

Find the inverse of the function f(x) = (x - 4)/(x + 4) x does not equal -4

OpenStudy (anonymous):

let y = (x-4)/(x+4) now switch the x's and the y's and solve for y. the expression you get is your inverse.

OpenStudy (anonymous):

So the answer would just be (x+4)/(x-4) = x^2+16. right?